1. a. Amy bought a baseball card for
$12. To date, the value of the
card increased by 30%, then
decreased by 15%, and finally
increased by 40%.
Joe bought a baseball card for
$12. To date, the value of the
card decreased by 10%, then
increased by 70%, and finally
decreased by 5%.
Whose card is currently worth
more? by how much? Explain.
b. By about what percent must the
currently lesser-valued card
increase to be of equal value with
the greater-valued card? Round
your answer to the nearest tenth
of a percent.
2. a. Jorge’s family bought a house for
$125,000. To date, the value of
the house increased by 5%, then
decreased by 25%, and finally
increased by 10%.
Gene’s family bought a house for
$125,000. To date, the value of
the house decreased by 5%, then
increased by 15%, and finally
decreased by 20%.
Whose house is currently worth
more? by how much? Explain.
b. By about what percent must the
currently lesser-valued house
increase to be of equal value with
the greater-valued house? Round
your answer to the nearest tenth
of a percent.
May you please show your work? Thank you, and have a great day!

Answer by addingup(3677)

12+(12*0.30) = 15.60
15.60-(15.60*0.15) = 13.26
13.26+(13.26*0.40) = your final answer.
Do the others the same way.